OpenStudy (anonymous):

Which is an equation of the line containing the points (4, 10) and (6, 11) in standard form? (Points : 2) –x + 2y = 16 2x – y = –2 x + 2y = 24 x – 2y = –16

OpenStudy (cwrw238):

first find the slope of the line by subtracting the values of the y coordinates and divide by the difference of x coordinates so we get (11-10) / (6-4) = 1/2 now use the general foemula y - y1 = m(x - x1) where (x1,y1) is some point on the line and m = slope

OpenStudy (cwrw238):

so plug in m = 1/2 , x1 = 4 and y1 = 10 into the above formula and simplfiy to get the answer

OpenStudy (ranga):

One way to answer this question is to try out the two points in each choice and see if it works. Put x = 4 and y = 10 and see if it works. Then try x = 6, y = 11 in each. The one that satisfies both is the correct answer. But if you want to know how to find the equations without trying out the answer choices let me know.

OpenStudy (cwrw238):

need more help?

OpenStudy (cwrw238):